# BSTcontrol

### v2

@@ 2.8 Deployment of Large AI Models in Wireless Networks
Deploying large AI models in wireless networks involves navigating complex trade-offs between computational efficiency, latency, privacy, and scalability. To meet the demands of real-time and resource-constrained environments, the feasibility of fine-tuning large AI models on an edge node in 6G networks primarily hinges on peak memory usage, which is dominated by model parameters, gradients, optimizer states, data batch size, intermediate activations, and fragmented memory from residual states [67].
@@ 2.8.1 Architectural Strategies and Optimization Techniques
Deploying large AI models in wireless communication systems requires architectural strategies that address constraints such as limited computation, stringent latency, and energy efficiency. Centralized deployments on high-performance GPU clusters that are typically hosted on public cloud platforms enable scalable inference for tasks like traffic prediction, network planning, and simulation. These platforms support serving large models via application programming interfaces (APIs) to multiple users or base stations. However, latency, bandwidth cost, and privacy concerns arise when transferring telemetry to remote servers. To mitigate these limitations, hybrid solutions can cache knowledge locally or offload selective tasks closer to the data source.
When deploying models closer to the network edge, compression techniques like pruning, quantization, and distillation become essential to reduce memory and compute demands [86, 87]. These optimizations allow models to run on smaller-scale hardware such as edge GPUs or embedded processors. Advanced deployment schemes such as federated learning enable local model training while preserving data privacy, and split learning approaches can distribute model inference between local and remote devices [88]. Such strategies are crucial for real-time applications like adaptive scheduling, beam tracking, and anomaly detection, which are critical for real-time applications like ultra-reliable low latency communications (URLLC). Adaptive mechanisms and continuous monitoring are also necessary to ensure robust performance under dynamic and unpredictable wireless conditions.
@@ 2.8.2 Hardware and Specialized Frameworks
Effective deployment also relies heavily on selecting suitable hardware and software frameworks. Inference engines must accommodate resource-constrained environments such as mobile devices, base stations, or multi-access edge computing (MEC) nodes. Specialized hardware accelerators like neural processing units enable efficient execution of AI models in these scenarios. On-device inference supports applications like personalized assistants, local RF analytics, and user-side diagnostics, offering benefits such as low-latency responses, offline functionality, and enhanced privacy [89].
To further support lightweight deployments, frameworks like LiteRT [90], PyTorch Mobile [91], and ONNX Runtime [92] provide infrastructure for running quantized or distilled models on constrained hardware. Efficient model architectures, such as minGPT [93], complement these frameworks by reducing computational complexity without compromising significantly on performance. In hybrid deployments, devices can dynamically offload intensive tasks to more capable nodes depending on real-time conditions, maintaining a balance between autonomy and performance. Security and ethical considerations must also be integrated throughout the deployment lifecycle to safeguard sensitive data and prevent model misuse. Altogether, these strategies enable large AI models to enhance wireless networks across domains such as signal processing, resource allocation, and intelligent network management.
@@ 2.9 Design Principles and Key Characteristics of WLAM Systems
Below, we highlight several critical aspects that must be considered to ensure the effectiveness and efficiency when designing methods that integrate large AI models with wireless communications. First, the adaptability of the models to dynamic environments is essential. Wireless networks experience fluctuating signal strengths, mobility, and interference. Thus, AI models must adjust in real-time without extensive retraining to maintain performance in unpredictable conditions. Given the constraints of wireless systems, such as latency, energy consumption, and computational limitations, AI models must be optimized for edge deployment. This involves designing models to function efficiently within the limited resources of edge devices, ensuring low energy consumption while maintaining performance. The integration of diverse data types is crucial. Wireless systems rely on radio-frequency signals, images, and text, so WLAM must seamlessly process and align these modalities to enable context-aware decision-making across different network conditions. Privacy and security concerns are critical. AI models in wireless systems must protect sensitive data, comply with regulations, and minimize biases, ensuring data security and privacy to maintain trust and meet legal requirements. Finally, WLAM should be adaptive. The models must adapt in real-time based on network feedback, ensuring continuous performance improvement without manual intervention. This feature is vital for supporting the evolving demands of next-generation wireless systems.
To sum up, integrating WLAM in wireless networks offers significant potential to address the growing complexity of communication systems. By effectively combining multiple resources, adapting to dynamic environments, and ensuring privacy and security, WLAM can enable more efficient, scalable, and intelligent wireless systems, essential for the success of 6G and beyond.
@@ 3 Large AI Models for Wireless Communications
Figure 6:   The outline of Section III.
The escalating complexity of modern wireless systems and the ever-increasing demand for enhanced performance present significant challenges. To overcome these challenges, large AI models are rapidly emerging as powerful tools within the field of wireless communications. Large AI models offer significant potential to address these challenges across various communication layers. This section delves into the application of large AI models across physical, network, and semantic layers, as well as their role as knowledge agents, as depicted in Fig. 6. The related works are summarized in Table 3. Subsequent subsections will detail how these models enhance the efficiency, robustness, and intelligence of wireless communications within each domain.
@@ 3.1 WLAM for Physical Layer Communications
Layer Ref. Scenarios Contributions
Physical Layer [94] Vision-aided beam prediction in mmWave vehicle-to-infrastructure communication systems using RGB images. Propose BeamLLM, leveraging LLMs and computer vision for high accuracy mmWave beam prediction.
[95] End-to-end channel coding for unknown/non-differentiable channels. Propose conditional diffusion models for channel decoding, showing a higher quality of generation in image-based tasks
[96] Multimodal ISAC systems, exemplified by beam prediction using multimodal sensor data (e.g., GPS, RGB images). Propose an MLLM-enabled framework for multimodal ISAC to enhance communication and sensing.
[97] Physical layer communication security optimization problems, specifically cooperative friendly jamming scenarios. Propose an MoE-enabled generative artificial intelligence framework to enhance security in cooperative jamming scenarios.
[99] Adapting LLMs for various networking tasks, including prediction and decision-making involving multimodal inputs. Propose NetLLM, the first framework to efficiently adapt LLMs for networking tasks.
[100] Providing personalized generative services in future AI-native networks via collaborative cloud-edge LLM deployment. Propose NetGPT, an AI-native network architecture synergizing cloud and edge LLMs for network management tasks.
Semantic Layer [101] Semantic wireless image transmission using JSCC leveraging deep generative models. Propose InverseJSCC and GenerativeJSCC leveraging StyleGAN-based generators for semantic image transmission.
[102] Wireless semantic communications over noisy channels, where channel noise degrades the received signal after equalization. Propose CDDM as a post-equalization module to remove channel noise by learning the input signal distribution.
[103] Semantic communication systems for text transmission over wireless channels, applying LLMs to encoding and decoding. Propose LLM-SC, the first framework using LLMs for physical layer semantic encoding and decoding.
Wireless Agents [104] Adapting general-purpose LLMs for telecom-specific tasks. Propose a framework and pipeline to create telecom-specific LLMs, showing strong performance on telecom tasks.
[105] FPGA-based hardware development for advanced wireless communication signal processing algorithms. Investigate LLM assistance in FPGA-based SDR development via a case study and identified key LLM uses.
[106] Agentic AI networking in 6G, supporting interaction, collaborative learning, and knowledge transfer among diverse AI agents. Propose AgentNet, a novel framework for agentic AI networking, and demonstrated its potential in two application scenarios.
[107] Automating the full lifecycle of complex 5G/6G network simulations using the ns-3 simulator. Propose an innovative multi-agent framework integrating LLMs with ns-3, and validate effectiveness through a 5G case study.
@@ 3.1.1 Channel-Associated Prediction
Channel-associated prediction tasks are a significant area within physical layer communications where large AI models show remarkable potential. Accurate prediction of CSI is crucial for adaptive resource management, beamforming optimization, and interference mitigation. Traditional channel prediction methods often struggle with rapid channel variations in mobile wireless environments and complex non-linear relationships inherent in wireless propagation. Large AI models, leveraging their powerful function approximation capabilities, can learn temporal and spatial channel dynamics from historical data and environmental context. For example, in beam prediction for millimeter wave systems, LLMs are being explored to forecast optimal beam directions using past beam indices and angles of departure [108]. Vision-aided techniques are also emerging. BeamLLM from [94] utilizes LLMs to process colored images for beam prediction, demonstrating high accuracy in vehicle-to-infrastructure scenarios. These models exhibit strong robustness and generalization capabilities, outperforming traditional learning based methods in prediction accuracy and adaptability to diverse wireless conditions.
@@
Enhancing the robustness of these predictions against the rapid channel variations and inherent uncertainties of wireless environments is important. Bayesian learning techniques, particularly Bayesian neural networks (BNNs), offer a principled framework for quantifying prediction uncertainty [109]. Unlike conventional models that provide point estimates, BNNs produce predictive distributions, distinguishing between aleatoric uncertainty (inherent data noise) and epistemic uncertainty (model uncertainty due to limited data). This uncertainty quantification enhances reliability, especially with limited training data or in the presence of outliers and model misspecification, common issues in wireless settings. While traditional BNN inference can be computationally intensive, approaches like real time teacher-student BNNs have been developed to enable fast uncertainty aware predictions suitable for dynamic applications like vehicle tracking [110]. Integrating such uncertainty awareness into WLAMs could significantly improve their robustness and decision making capabilities under fluctuating channel conditions.
@@
Furthermore, the accuracy and robustness of channel predictions can be substantially improved by fusing information from diverse data sources. Wireless systems generate heterogeneous data streams, including CSI, beam training logs, and environmental sensor data such as camera images, LiDAR point clouds, or radar signals. WLAM, often built upon transformer architectures, inherently possesses powerful mechanisms like attention for multimodal data fusion [111]. Attention based fusion allows the model to dynamically weigh the importance of different data sources or features based on the current context, effectively integrating complementary information and enhancing prediction accuracy. For example, visual data can provide context for blockages, while CSI captures temporal dynamics. To further enhance robustness in noisy environments, models can be designed to account for data quality, or Bayesian uncertainty estimates can be used to give less weight to unreliable sources.
@@
Building upon these advancements in specific prediction tasks, general frameworks are being developed to harness large AI models for a broader range of wireless channel-associated problems. The LLM4WM framework, detailed in [112], introduces a comprehensive approach for adapting LLMs to these challenges. This framework employs a MoE with LoRA to achieve efficient multi task fine tuning, thus enabling the transfer of pre-trained LLM knowledge to diverse wireless communication scenarios. Furthermore, the effectiveness of fine-tuned LLMs for general channel prediction is highlighted by LLM4CP, presented in [85], which achieves state of the art performance in both time division duplex and frequency division duplex systems. By leveraging large AI models and general frameworks like LLM4WM for channel-associated prediction tasks, wireless systems can proactively adapt to channel fluctuations, optimize transmission parameters, and enhance overall network performance across multiple functionalities.

### v4

@@ 2.8 Deployment of Large AI Models in Wireless Networks
Deploying large AI models in wireless networks involves navigating complex trade-offs between computational efficiency, latency, privacy, and scalability. To meet the demands of real-time and resource-constrained environments, the feasibility of fine-tuning large AI models on an edge node in 6G networks primarily hinges on peak memory usage, which is dominated by model parameters, gradients, optimizer states, data batch size, intermediate activations, and fragmented memory from residual states [yu2024snake].
@@ 2.8.1 Architectural Strategies and Optimization Techniques
Deploying large AI models in wireless communication systems requires architectural strategies that address constraints such as limited computation, stringent latency, and energy efficiency. Centralized deployments on high-performance GPU clusters that are typically hosted on public cloud platforms enable scalable inference for tasks like traffic prediction, network planning, and simulation. These platforms support serving large models via APIs to multiple users or base stations. However, latency, bandwidth cost, and privacy concerns arise when transferring telemetry to remote servers. To mitigate these limitations, hybrid solutions can cache knowledge locally or offload selective tasks closer to the data source.
When deploying models closer to the network edge, compression techniques like pruning, quantization, and distillation become essential to reduce memory and compute demands [han2015deep, xiao2023smoothquant]. These optimizations allow models to run on smaller-scale hardware such as edge GPUs or embedded processors. Advanced deployment schemes such as federated learning enable local model training while preserving data privacy, and split learning approaches can distribute model inference between local and remote devices [kang2017neurosurgeon]. Such strategies are crucial for real-time applications like adaptive scheduling, beam tracking, and anomaly detection, which are critical for real-time applications like ultra-reliable low latency communications (URLLC). Adaptive mechanisms and continuous monitoring are also necessary to ensure robust performance under dynamic and unpredictable wireless conditions.
@@ 2.8.2 Hardware and Specialized Frameworks
Effective deployment also relies heavily on selecting suitable hardware and software frameworks. Inference engines must accommodate resource-constrained environments such as mobile devices, base stations, or multi-access edge computing (MEC) nodes. Specialized hardware accelerators like neural processing units enable efficient execution of AI models in these scenarios. On-device inference supports applications like personalized assistants, local RF analytics, and user-side diagnostics, offering benefits such as low-latency responses, offline functionality, and enhanced privacy [deng2020edge].
To further support lightweight deployments, frameworks like LiteRT [litert], PyTorch Mobile [pytorchmobile], and ONNX Runtime [onnxrun] provide infrastructure for running quantized or distilled models on constrained hardware. Efficient model architectures, such as minGPT [mingpt], complement these frameworks by reducing computational complexity without compromising significantly on performance. In hybrid deployments, devices can dynamically offload intensive tasks to more capable nodes depending on real-time conditions, maintaining a balance between autonomy and performance. Security and ethical considerations must also be integrated throughout the deployment lifecycle to safeguard sensitive data and prevent model misuse. Altogether, these strategies enable large AI models to enhance wireless networks across domains such as signal processing, resource allocation, and intelligent network management.
Table 4: Architectural Comparison: 5G + AI vs. AI-Native 6G.
Dimension 5G + AI AI-Native 6G (WLAM-based)
Integration Depth AI optimizes specific tasks like channel estimation and beamforming within existing telecommunication protocols. The protocol stack and air interface are designed with native AI integration, such as semantic communications.
Model Paradigm Task-specific small models are employed, which requires retraining for new scenarios. Foundation models are trained for all tasks, enabling general knowledge transfer and few-shot learning.
Data Processing Data processing is managed at its own dedicated layer, such as separate source and channel coding. Multimodal data including RF, vision, and text are fused and processed in an end-to-end mode.
Deployment Strategy Models generally operate within the individual network devices while utilizing standard communication links for any necessary coordination. The network architecture natively supports distributed deployment mechanisms such as split computing to accommodate massive models on constrained edge devices.
Operation Mode Tasks are executed by models based on strictly predefined algorithms, requiring constant human intervention and monitoring. Agents autonomously perceive the environment, reason about their goals, formulate complex plans, and execute intricate workflows.
@@ 2.9 Design Principles and Key Characteristics of WLAM Systems
Below, we highlight several critical aspects that must be considered to ensure the effectiveness and efficiency when designing methods that integrate large AI models with wireless communications. First, the adaptability of the models to dynamic environments is essential. Wireless networks experience fluctuating signal strengths, mobility, and interference. Thus, AI models must adjust in real-time without extensive retraining to maintain performance in unpredictable conditions. Given the constraints of wireless systems, such as latency, energy consumption, and computational limitations, WLAMs must be optimized for edge deployment. This involves designing models to function within the limited resources of edge devices while maintaining performance. The integration of diverse data types is crucial. Wireless systems rely on radio-frequency signals, images, and text, so WLAM must seamlessly process and align these modalities to enable context-aware decision-making across varying network conditions. Privacy and security concerns are critical. AI models in wireless systems must protect sensitive data, comply with regulations, and minimize biases, ensuring data security and privacy to maintain trust and meet legal requirements. Finally, WLAM should be adaptive. The models must adapt in real-time based on network feedback, ensuring continuous performance improvement without manual intervention. This feature is vital for supporting the evolving demands of next-generation wireless systems.
To further clarify the transformative role of WLAMs within these design principles, it is essential to distinguish the vision of an AI-Native 6G system from the fifth generation of cellular network technology (5G) systems enhanced with AI capabilities. The fundamental difference lies not merely in the model size, but in the architectural philosophy and operational paradigm. In 5G, AI is typically applied as an add-on feature to optimize specific, isolated modules (e.g., AI-based CSI feedback or beam management) within a fixed protocol framework. This represents a “one model for one task” paradigm, where models lack cross-task generalization. In contrast, an AI-Native 6G system implies that the air interface and network architecture are designed from scratch with WLAMs at their core [fontaine2024towards]. This shift moves towards “one foundation model for all tasks,” enabling the network to handle unseen scenarios through the inherent generalization and reasoning capabilities of WLAMs [wirelessgpt, LLM4WM]. Simultaneously, regarding data processing, the paradigm evolves from the separate source and channel coding blocks of 5G to an end-to-end approach, where WLAMs intrinsically fuse and process multimodal data including RF signals, vision, and text [cheng2025large]. In terms of deployment, the system transitions from operating within the individual nodes to a natively collaborative edge-cloud synergy, where techniques like split computing are essential to accommodate massive model parameters. Furthermore, AI-Native 6G introduces agentic AI, where the system possesses the autonomy to decompose high-level intents into executable network configurations [jiang2025largeaimodelsagentic]. Table 4 provides a structured comparison of these criteria.
To sum up, integrating WLAM in wireless networks offers significant potential to address the growing complexity of communication systems. Specifically, the diverse enabling technologies discussed in this section map directly to critical 6G performance goals and deployment constraints: advanced neural architectures like Transformers and scaling strategies such as MoE enable the high-capacity representation learning necessary for strong generalization across diverse wireless scenarios; multimodal learning bridges heterogeneous data modalities to enhance context-awareness and sensing precision; diffusion-based models leverage iterative denoising processes to enhance robustness against non-linear physical layer distortions; reinforcement learning frameworks empower the network with dynamic adaptability and high-level autonomy; parameter-efficient fine-tuning addresses the stringent energy and storage efficiency constraints of edge deployment; and advanced inference strategies such as RAG and CoT significantly improve the reliability and factual accuracy of agentic decisions. By effectively combining multiple resources, adapting to dynamic environments, and ensuring privacy and security, WLAM can enable more efficient, scalable, and intelligent wireless systems essential for the success of 6G and beyond. We provide a summarization of the state-of-the-art technologies regarding these design principles in Table 3.
@@ 3 Large AI Models for Wireless Communications
Figure 7:   The outline of Section 3.
The escalating complexity of modern wireless systems and the ever-increasing demand for enhanced performance present significant challenges. To overcome these, large AI models are rapidly emerging as powerful tools within the field of wireless communications. Building on the fundamentals established in Section 2, the key to this transformative potential lies in the exceptional generalization and reasoning capabilities of WLAMs. Unlike traditional models that struggle with unseen channel distributions, the powerful generalization capability of WLAMs derived from massive-scale pre-training forms the cornerstone for adapting to the stochastic, time-varying wireless environments. With this logical foundation, this section delves into the specific applications of WLAMs by systematically progressing through the communication stack, as depicted in Fig. 7. Our discussion begins at the physical layer, moves to the network and semantic layers, and explores the emerging concept of wireless agents. Finally, these diverse applications are synthesized into a structured framework to provide practitioners with practical implementation guidelines. An overview of the related works is provided in Table 5.
Layer Ref. Scenarios Contributions
Physical Layer [LLM4CP] Adapting LLMs for high-performance wireless channel prediction in TDD and FDD systems. Propose LLM4CP, demonstrating that a fine-tuned LLM can achieve state-of-the-art performance for wireless channel prediction.
[LLM4WM] Adapting a single pre-trained LLM to multiple, diverse wireless communication tasks. Propose LLM4WM, a general framework for efficient multi-task fine-tuning of LLMs in various wireless scenarios.
[10104549] End-to-end channel coding for unknown/non-differentiable channels. Leverage generative diffusion models, a key architecture in modern large AI, for end-to-end channel decoding.
[cheng2025large] Multimodal ISAC systems, exemplified by beam prediction using multimodal sensor data (e.g., GPS, RGB images). Propose an MLLM-enabled framework for multimodal ISAC to enhance communication and sensing.
[wcmsecurity] Physical layer communication security optimization problems, specifically cooperative friendly jamming scenarios. Propose an MoE-enabled generative artificial intelligence framework to enhance security in cooperative jamming scenarios.
[NetLLM] Adapting LLMs for various networking tasks, including prediction and decision-making involving multimodal inputs. Propose NetLLM, the first framework to efficiently adapt LLMs for networking tasks.
[chen2024netgpt] Providing personalized generative services in future AI-native networks via collaborative cloud-edge LLM deployment. Propose NetGPT, an AI-native network architecture synergizing cloud and edge LLMs for network management tasks.
Semantic Layer [wu2024cddm] Wireless semantic communications over noisy channels, where channel noise degrades the received signal after equalization. Propose CDDM as a post-equalization module to remove channel noise by learning the input signal distribution.
[wang2024large] Semantic communication systems for text transmission over wireless channels, applying LLMs to encoding and decoding. Propose LLM-SC, the first framework using LLMs for physical layer semantic encoding and decoding.
[ni2024interplay] Exploring the synergistic interplay between semantic communication and the knowledge base of large AI models. Investigate the synergistic interplay where knowledge within a large model is leveraged for communication efficiency.
Wireless Agents [zou2024telecomgpt] Adapting general-purpose LLMs for telecom-specific tasks. Propose a framework and pipeline to create telecom-specific LLMs, showing strong performance on telecom tasks.
[du2024powerlargelanguagemodels] FPGA-based hardware development for advanced wireless communication signal processing algorithms. Investigate LLM assistance in FPGA-based SDR development via a case study and identified key LLM uses.
[xiao2025agenticainetworking6g] Agentic AI networking in 6G, supporting interaction, collaborative learning, and knowledge transfer among diverse AI agents. Propose AgentNet, a novel framework for agentic AI networking, and demonstrated its potential in two application scenarios.
[rezazadeh2025generative6gsimulationexperimental] Automating the full lifecycle of complex 5G/6G network simulations using the ns-3 simulator. Propose an innovative multi-agent framework integrating LLMs with ns-3, and validate effectiveness through a 5G case study.
@@ 3.1.1 Channel-Associated Prediction
Channel-associated prediction tasks are a significant area within physical layer communications where large AI models show remarkable potential. Accurate prediction of CSI is crucial for adaptive resource management, beamforming optimization, and interference mitigation. Traditional channel prediction methods often struggle with rapid channel variations in mobile wireless environments and complex non-linear relationships inherent in wireless propagation. Large AI models, leveraging their powerful function approximation capabilities, can learn temporal and spatial channel dynamics from historical data and environmental context. For example, in beam prediction for millimeter wave systems, LLMs are being explored to forecast optimal beam directions using past beam indices and angles of departure [Sheng2025]. Vision-aided techniques are also emerging. BeamLLM from [zheng2025beamllmvisionempoweredmmwavebeam] utilizes LLMs to process colored images for beam prediction, demonstrating high accuracy in vehicle-to-infrastructure scenarios. These models exhibit strong robustness and generalization capabilities, outperforming traditional learning based methods in prediction accuracy and adaptability to diverse wireless conditions.
@@
Enhancing the robustness of these predictions against the rapid channel variations and inherent uncertainties of wireless environments is important. Bayesian learning techniques, particularly Bayesian neural networks (BNNs), offer a principled framework for quantifying prediction uncertainty [bayesianTCCN]. Unlike conventional models that provide point estimates, BNNs produce predictive distributions, distinguishing between aleatoric uncertainty (inherent data noise) and epistemic uncertainty (model uncertainty due to limited data). This uncertainty quantification enhances reliability, especially with limited training data or in the presence of outliers and model misspecification, common issues in wireless settings. While traditional BNN inference can be computationally intensive, approaches like real time teacher-student BNNs have been developed to enable fast uncertainty aware predictions suitable for dynamic applications like vehicle tracking [bayesianJSAC]. Integrating such uncertainty awareness into WLAMs could significantly improve their robustness and decision making capabilities under fluctuating channel conditions.
@@
Furthermore, the accuracy and robustness of channel predictions can be substantially improved by fusing information from diverse data sources. Wireless systems generate heterogeneous data streams, including CSI, beam training logs, and environmental sensor data such as camera images, LiDAR point clouds, or radar signals. WLAM, often built upon transformer architectures, inherently possesses powerful mechanisms like attention for multimodal data fusion [ZHOU2023126604]. Attention based fusion allows the model to dynamically weigh the importance of different data sources or features based on the current context, effectively integrating complementary information and enhancing prediction accuracy. For example, visual data can provide context for blockages, while CSI captures temporal dynamics. To further enhance robustness in noisy environments, models can be designed to account for data quality, or Bayesian uncertainty estimates can be used to give less weight to unreliable sources.
@@
Building upon these advancements in specific prediction tasks, general frameworks are being developed to harness large AI models for a broader range of wireless channel-associated problems. The LLM4WM framework, detailed in [LLM4WM], introduces a comprehensive approach for adapting LLMs to these challenges. This framework employs a MoE with LoRA to achieve efficient multi task fine tuning, thus enabling the transfer of pre-trained LLM knowledge to diverse wireless communication scenarios. Furthermore, the effectiveness of fine-tuned LLMs for general channel prediction is highlighted by LLM4CP, presented in [LLM4CP], which achieves state of the art performance in both time division duplex and frequency division duplex systems. By leveraging large AI models and general frameworks like LLM4WM for channel-associated prediction tasks, wireless systems can proactively adapt to channel fluctuations, optimize transmission parameters, and enhance overall network performance across multiple functionalities.